OpenStudy (anonymous):

Can Someone Solve these & Show The Work Please Solve the systems using substitution. 6x + 2y = 0 4x + y = -1 3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18

OpenStudy (anonymous):

to start off substitution 2 variables solve for one variable first

OpenStudy (anonymous):

EXAMPLE X2 + Y2 = 2 SOLVE FOR X USING THIS EQUATION X3 + Y3 = 3 1/2 [ X2 + Y2 = 2 ] X + Y = 2 X = - Y + 2 THIS IS WHAT X IS EQUAL TO X3 + Y3 = 3 USE THE NEXT EQUATION 1/3 [ X3 + Y3 = 3 ] X + Y = 1 ( -Y +2 ) + Y = = 1 |__________| THIS IS SUBSTITUTED AS X - Y + Y = 1 -2 0 = - 2 NO REAL NUMBER SOL. BUT THE PROCESS IS THE SAME SUBSTITUTION WITH 2 VARIABLES